OpenStudy (anonymous):

Solve the equation. 2(y + 7) = 4y + 14

OpenStudy (anonymous):

2(y+7) = 4y + 14 2y + (2)(7) = 4y +14 14 = 2y + 14 0 = 2y 0 = y

OpenStudy (anonymous):

2y-4y=0 y=0